INTERMEDIATE OPTIMAL GEVREY EXPONENTS OCCUR

MICHAEL CHRIST

Abstract. Hypoellipticity in Gevrey classes G^{s} is characterized for a simple family of sums of squares of vector fields satisfying the bracket hypothesis, with analytic coefficients. It is shown that hypoellipticity holds if and only if s is greater than or equal to an optimal exponent that may take on any rational value.